# Condensed-space methods for nonlinear programming on GPUs

### v1

@@ 3.1 Golub & Greif strategy: HyKKT
The Golub & Greif golub2003solving strategy reformulates the KKT system using an Augmented Lagrangian formulation. It has been recently revisited in regev2023hykkt to solve the condensed KKT system ($K_{1}$) on the GPU. For a dual regularization $\delta_{c}=0$, the trick is to reformulate the condensed KKT system ($K_{1}$) in an equivalent form
@@
where we have introduced the regularized matrix $K_{\gamma}:=K+\gamma G^{\top}G$. We note by $Z$ a basis of the null-space of the Jacobian $G$. Using a classical result from debreu1952definite , if $G$ is full row-rank then there exists a threshold value $\underline{\gamma}$ such that for all $\gamma>\underline{\gamma}$, the reduced Hessian $Z^{\top}KZ$ is positive definite if and only if $K_{\gamma}$ is positive definite. Using the Sylvester’s law of inertia stated in (9), we deduce that for $\gamma>\underline{\gamma}$, if $\operatorname{inertia}(K_{2})=(n+m_{i},m_{e}+m_{i},0)$ then $K_{\gamma}$ is positive definite.
The linear solver HyKKT regev2023hykkt leverages the positive definiteness of $K_{\gamma}$ to solve (10) using a hybrid direct-iterative method that uses the following steps:
@@
The method uses a sparse Cholesky factorization along with the conjugate gradient (Cg) algorithm hestenes-stiefel-1952 . The sparse Cholesky factorization has the advantage of being stable without numerical pivoting, rendering the algorithm tractable on a GPU. Each Cg iteration requires the application of sparse triangular solves with the factors of $K_{\gamma}$. For that reason, HyKKT is efficient only if the Cg solver converges in a small number of iterations. Fortunately, the eigenvalues of the Schur-complement $S_{\gamma}:=GK_{\gamma}^{-1}G^{\top}$ all converge to $\frac{1}{\gamma}$ as we increase the regularization parameter $\gamma$ (regev2023hykkt, , Theorem 4), implying that $\lim_{\gamma\to\infty}\kappa_{2}(S_{\gamma})=1$. Because the convergence of the Cg method depends on the number of distinct eigenvalues of $S_{\gamma}$, the larger the $\gamma$, the faster the convergence of the Cg algorithm in (11). Cr hestenes-stiefel-1952 and cAr montoison-orban-saunders-2023 can also be used as an alternative to Cg. Although we observe similar performance, these methods ensure a monotonic decrease in the residual norm of (11) at each iteration.
@@ 3.2 Equality relaxation strategy: LiftedKKT
The problem (12) has only inequality constraints. After introducing slack variables, the condensed KKT system ($K_{1}$) reduces to
@@
with $H_{\tau}=\big{(}G^{\top}~{}H^{\top}\big{)}^{\top}$ and $K_{\tau}:=W+D_{x}+\delta_{w}I+H_{\tau}^{\top}D_{H}H_{\tau}$. Using the relation (9), the matrix $K_{\tau}$ is guaranteed to be positive definite if the primal regularization parameter $\delta_{w}$ is adequately large. As such, the parameter $\delta_{w}$ is chosen dynamically using the inertia information of the system in ($K_{1}$). Therefore, $K_{\tau}$ can be factorized with a Cholesky decomposition, satisfying the key requirement of stable pivoting for the implementation on the GPU. The relaxation causes error in the final solution. Fortunately, the error is in the same order of the solver tolerance, thus it does not significantly deteriorate the solution quality for small $\varepsilon_{tol}$.
While this method can be implemented with small modification in the optimization solver, the presence of tight inequality in (12) causes severe ill-conditioning throughout the IPM iterations. Thus, using an accurate iterative refinement algorithm is necessary to get a reliable convergence behavior.
@@ 3.3 Discussion
We have introduced two algorithms to solve KKT systems on the GPU. On the contrary to classical implementations, the two methods do not require computing a sparse $\mathrm{LBL^{T}}$ factorization of the KKT system and use instead alternate reformulations based on the condensed KKT system ($K_{1}$). Both strategies rely on a Cholesky factorization: HyKKT factorizes a positive definite matrix $K_{\gamma}$ obtained with an Augmented Lagrangian strategy whereas Lifted KKT factorizes a positive definite matrix $K_{\tau}$ after using an equality relaxation strategy. We will see in the next section that the ill-conditioned matrices $K_{\gamma}$ and $K_{\tau}$ have a specific structure that limits the loss of accuracy in IPM.
@@ 4 Conditioning of the condensed KKT system
The relative errors are bounded above by a term depending on the conditioning $\kappa_{2}(M)$. Hence, it is legitimate to investigate the impact of the ill-conditioning when solving the condensed system ($K_{1}$) with LiftedKKT or with HyKKT. We will see that we can tighten the bounds in (14) by exploiting the structured ill-conditioning of the condensed matrix $K$. We base our analysis on wright1998ill , where the author has put a particular emphasis on the condensed KKT system ($K_{1}$) without equality constraints. We generalize her results to the matrix $K_{\gamma}$, which incorporates both equality and inequality constraints. The results extend directly to $K_{\tau}$ (by letting the number of equalities to be zero).
To ease the notation, we suppose that the primal variable $x$ is unconstrained, leaving only a bounded slack variable $s$ in (2). This is equivalent to include the bounds on the variable $x$ in the inequality constraints, as $\bar{h}(x)\leq 0$ with $\bar{h}(x):=(h(x),-x)$.
@@ 4.1 Centrality conditions
We start the discussion by recalling important results porting on the iterates of the interior-point algorithm. Let $p=(x,s,y,z)$ the current primal-dual iterate (the bound multiplier $v$ is considered apart), and $p^{\star}$ a solution of the KKT conditions (4). We suppose that the solution $p^{\star}$ is regular and satisfies the following assumptions. We denote $\mathcal{B}=\mathcal{B}(x^{\star})$ the active-set at the optimal solution $x^{\star}$, and $\mathcal{N}=\mathcal{N}(x^{\star})$ the inactive set.
@@
Let $p^{\star}=(x^{\star},s^{\star},y^{\star},z^{\star})$ be a primal-dual solution satisfying the KKT conditions (4). Let the following hold:
@@
$\delta(p,v)=\Theta\left(\left\|\begin{bmatrix}\nabla_{p}L(p,v)\\ \min(v,s)\end{bmatrix}\right\|\right)\;.$ (15)
@@
for some constants $C>0$ and $\alpha\in(0,1)$. Conditions (17b) ensures that the products $s_{i}v_{i}$ are not too disparate in the diagonal term $D_{s}$. This condition is satisfied (despite rather loosely) in the solver Ipopt (see (wachter2006implementation, , Equation (16))).
@@ 4.2.1 Invariant subspaces in $K_{\gamma}$
We remind that $m_{e}$ is the number of equality constraints, and define $\ell:=m_{e}+m_{a}$.
We explicit the structured ill-conditioning of $K_{\gamma}$ by modifying the approach outlined in (wright1998ill, , Theorem 3.2) to account for the additional term $\gamma G^{\top}G$ arising from the equality constraints. We show that the matrix $K_{\gamma}$ has two invariant subspaces (in the sense defined in (stewart1990matrix, , Chapter 5)), associated respectively to the range of the transposed active Jacobian (large space) and to the null space of the active Jacobian (small space).
@@
    In late IPM iterations, the active slacks are converging to 00. We observe that if $\frac{1}{\Xi}\leq\gamma\leq\frac{1}{s_{min}}$ the parameter $\gamma$ does not increase the ill-conditioning of the condensed matrix $K_{\gamma}$.
@@ 4.2.3 Numerical solution of the condensed system
Condition (a) implies that $s_{min}=\Theta(\Xi)$ and $s_{max}=\Theta(\Xi)$ (Proposition 1). Condition (b) supposes in addition $\gamma=\Theta(\Xi^{-1})$, making the matrix $\Sigma_{L}$ well-conditioned with $\underline{\sigma}=\Theta(\Xi^{-1})$, $\overline{\sigma}=\Theta(\Xi^{-1})$ and $\overline{\sigma}/\underline{\sigma}=\Theta(1)$. Condition (c) ensures that $\kappa_{2}(K_{\gamma})=\Theta(\overline{\sigma})$ satisfies $\kappa_{2}(K_{\gamma})\mathbf{u}\ll 1$ (implying the Cholesky factorization runs to completion). Condition (d) tells us that the perturbation caused by the Cholesky factorization is $\Delta_{s}K_{\gamma}=O(\mathbf{u}\|K_{\gamma}\|)$. As (28) implies $\|K_{\gamma}\|=\Theta(\Xi^{-1})$, we can incorporate $\Delta_{s}K_{\gamma}$ in the perturbation $\Delta K_{\gamma}$ given in Proposition 2.
@@ 4.3 Solution of the condensed KKT system
[…]
@@ 4.3.3 Summary
Most of the error arises in the active components $\widehat{d}_{z,\mathcal{B}}$. To limit the loss of accuracy, the algorithm has to decrease the absolute precision $\varepsilon_{K}$ as we are approaching to a local optimum. The impact remains limited if we have only few active inequality constraints.
5 Numerical results
We have implemented LiftedKKT and HyKKT on the GPU. First, we detail in §5.1 our implementation and assess in §5.2 the performance on the GPU of the two hybrid solvers Then, we present in §5.3 the results reported on the PGLIB OPF benchmark, complemented in §5.4 by the COPS benchmark.
5.1 Implementation
All our implementation uses the Julia language bezanson-edelman-karpinski-shah-2017 . We have used our local workstation to generate the results on the CPU, here equipped with an AMD EPYC 7443 processor (3.2GHz). For the results on the GPU, we have used an NVIDIA A100 GPU (with CUDA 12.3) on the Polaris testbed at Argonne National Laboratory  [^1] https://www.alcf.anl.gov/polaris.
@@ 5.1 Implementation
We have implemented the two condensed-space methods in our nonlinear IPM solver MadNLP shin2021graph . This implementation utilizes the abstraction AbstractKKTSystem in MadNLP to represent the various formulations of the KKT linear systems. MadNLP can deport most of the IPM algorithm to the GPU, except for basic IPM operations used for the coordination (e.g., the filter line-search algorithm). In particular, any operation that involves the manipulation of array entries is performed by GPU kernels without transferring data to host memory. We refer to shin2023accelerating for a detailed description of the GPU implementation in MadNLP.
@@
We use the ExaModels.jl modeling tool shin2023accelerating to implement all the nonlinear programs utilized in our benchmark. ExaModels.jl harnesses the sparsity structure and provides custom derivative kernels for repetitive algebraic subexpressions of the constraints and objective functions to compute first and second-order derivatives on the GPU in parallel bischof1991exploiting ; enzyme2021 . This approach caters to the SIMD architecture of the GPU by assigning each expression to multiple threads responsible for computing derivatives for different values.
@@
We solve the KKT systems assembled within MadNLP using various sparse linear solvers, chosen based on the KKT formulation ($K_{1}$, $K_{2}$, $K_{3}$) and the device (CPU, GPU) being utilized. We utilize the following solvers:
  • HSL MA27/MA57: Implement the $\mathrm{LBL^{T}}$ factorization on the CPU duff1983multifrontal . It solves the augmented KKT system ($K_{2}$). This solver serves as the reference when running on the CPU.
  • CHOLMOD: Implements the Cholesky and $\mathrm{LDL^{T}}$ factorizations on the CPU (using the AMD ordering amestoy-david-duff-2004 by default). It factorizes the condensed matrices $K_{\gamma}$ and $K_{\tau}$ appearing resp. in (10) and in (13). This solver is used to assess the performance of the hybrid solvers when running on the CPU.
  • cuDSS: Implement Cholesky, $\mathrm{LDL^{T}}$ and $\mathrm{LU}$ decompositions on an NVIDIA GPU. We use the $\mathrm{LDL^{T}}$ factorization to factorize the ill-conditioned condensed matrices $K_{\gamma}$ on GPU, as $\mathrm{LDL^{T}}$ is more robust than the Cholesky factorization.
@@
5.2 Performance analysis on a large-scale instance
We evaluate the performance of each KKT solver on a large-scale OPF instance, taken from the PGLIB benchmark babaeinejadsarookolaee2019power : 78484epigrids. Our formulation with ExaModels has a total of 674,562 variables, 661,017 equality constraints and 378,045 inequality constraints. Our previous work has pointed out that as soon as the OPF model is evaluated on the GPU using ExaModels, the KKT solver becomes the bottleneck in the numerical implementation shin2023accelerating .
5.2.1 Individual performance of the linear solvers
Subsequently, we evaluate the individual performance of the cuDSS solver when factorizing the matrix $K_{\gamma}$ at the first IPM iteration (here with $\gamma=10^{7}$). We compare the times to perform the symbolic analysis, the factorization and the triangular solves with those reported in CHOLMOD.
The results are displayed in Table 1. We benchmark the three decompositions implemented in cuDSS (Cholesky, $\mathrm{LDL^{T}}$, $\mathrm{LU}$), and assess the absolve accuracy of the solution by computing the infinity norm of the residual $\|K_{\gamma}x-b\|_{\infty}$. We observe that the analysis phase is four times slower for cuDSS compared to CHOLMOD. Fortunately, this operation needs to be computed only once in IPM, meaning that its cost is amortized if we run many IPM iterations. The factorization is about twenty times faster in cuDSS, with a time almost independent of the algorithm being used. The backward and forward sweeps are ten times faster: the triangular solves are harder to parallelize on a GPU. In terms of accuracy, the quality of the solution remains on par with CHOLMOD, except for the $\mathrm{LDL^{T}}$ decomposition which lags behind by at least two orders of magnitude.
linear solver analysis (s) factorization (s) backsolve (s) abs. accuracy
CHOLMOD 1.18 $8.57\times 10^{-1}$ $1.27\times 10^{-1}$ $3.60\times 10^{-13}$
cuDSS-Cholesky 4.52 $3.75\times 10^{-2}$ $1.32\times 10^{-2}$ $2.64\times 10^{-13}$
cuDSS-$\mathrm{LU}$ 4.50 $3.72\times 10^{-2}$ $1.49\times 10^{-2}$ $2.58\times 10^{-13}$
cuDSS-$\mathrm{LDL^{T}}$ 4.50 $4.07\times 10^{-2}$ $1.55\times 10^{-2}$ $7.62\times 10^{-11}$
Table 1: Comparing the performance of cuDSS with CHOLMOD. The matrix $K_{\gamma}$ is symmetric positive definite, with a size $n=674,562$. The matrix is extremely sparse, with only $7,342,680$ non-zero entries ($0.002$%).

### v2

@@ 3.1 Golub & Greif strategy: HyKKT
The Golub & Greif golub2003solving strategy reformulates the KKT system using an Augmented Lagrangian formulation. It has been recently revisited in regev2023hykkt to solve the condensed KKT system ($K_{1}$) on the GPU. The key idea is to reformulate the condensed KKT system ($K_{1}$) in an equivalent form:
@@
where we introduce the regularized matrix $K_{\gamma}:=K+\gamma G^{\top}G$ (recall $\delta_{c}=0$ is assumed per Remark 1) . Let $\begin{bmatrix}Y&Z\end{bmatrix}$ be an orthogonal matrix, where $Z$ encodes the basis of the null-space of the Jacobian $G$. Using a classical result from debreu1952definite , if $G$ is full row-rank, then there exists a threshold value $\underline{\gamma}$ such that for all $\gamma>\underline{\gamma}$, the reduced Hessian $Z^{\top}KZ$ is positive definite if and only if $K_{\gamma}$ is positive definite. Using Sylvester’s law of inertia stated in (9), we deduce that for $\gamma>\underline{\gamma}$, $\operatorname{inertia}(K_{2})=(n+m_{i},m_{e}+m_{i},0)$ if and only if the condensed matrix $K_{\gamma}$ is positive definite.
The linear solver HyKKT regev2023hykkt leverages the positive definiteness of $K_{\gamma}$ to solve (11) using a hybrid direct-iterative method with the following steps:
@@
The method employs a sparse Cholesky factorization along with the conjugate gradient (Cg) algorithm hestenes-stiefel-1952 . The sparse Cholesky factorization is advantageous as it is stable without numerical pivoting, making the algorithm tractable on a GPU. Each Cg iteration requires applying sparse triangular solves with the factors of $K_{\gamma}$. For this reason, HyKKT is efficient only if the Cg solver converges in few iterations. Fortunately, the eigenvalues of the Schur complement $S_{\gamma}:=GK_{\gamma}^{-1}G^{\top}$ all converge to $\frac{1}{\gamma}$ as we increase the regularization parameter $\gamma$ (regev2023hykkt, , Theorem 4), implying that $\lim_{\gamma\to\infty}\kappa_{2}(S_{\gamma})=1$. Since the convergence of the Cg method depends on the number of distinct eigenvalues of $S_{\gamma}$, the larger the $\gamma$, the faster the convergence of the Cg algorithm in (12). Although we observe similar performance, these methods ensure a monotonic decrease in the residual norm of (12) at each iteration.
Remark 2
The parameter $\gamma$ should be set above a given threshold $\underline{\gamma}$ for $K_{\gamma}$ to ensure positive definiteness of the system in (12). However, from the numerical stand-point, it is desired to choose moderate values of $\gamma$, since large $\gamma$ increases the ill-conditioning in $K_{\gamma}$, hence deteriorating the quality of the solution. In golub2003solving , the authors recommend choosing $\gamma=\|K\|/\|G\|^{2}$, which offers a good empirical trade-off. In (gould2010spectral, , Lemma 2.10), a tighter estimate is provided for the threshold $\underline{\gamma}$ if $G$ has full row-rank and $Z^{\top}KZ$ is positive definite:
$\underline{\gamma}=\dfrac{1}{\sigma_{\min}^{2}}\left(\dfrac{\|K\|^{2}}{\lambda_{\min}^{Z}}-\lambda_{\min}^{Y}\right)\;,$ (13)
with $\sigma_{\min}$ being the smallest nonzero singular value of the Jacobian $G$, $\lambda_{\min}^{Z}>0$ the smallest eigenvalue of $Z^{\top}KZ$, and $\lambda_{\min}^{Y}$ the smallest eigenvalue of $Y^{\top}KY$. From (13), we observe that $\underline{\gamma}$ is inversely proportional to $\lambda_{\min}^{Z}$ and $\sigma_{\min}$, indicating that it can reach a large magnitude if $G$ has nearly aligned rows or if $Z^{\top}KZ$ is close to being indefinite. In other words, HyKKT may have difficulties in solving problems with nearly dependent equality constraints or when the reduced Hessian is close to being indefinite.
@@ 3.2 Equality Relaxation Strategy: LiftedKKT
The problem (14) has only inequality constraints. As a result, the condensed KKT system ($K_{1}$) reduces to
@@
with $H_{\tau}=\big{(}G^{\top}~H^{\top}\big{)}^{\top}$ and $K_{\tau}:=W+\delta_{w}I+H_{\tau}^{\top}D_{H}H_{\tau}$ (recall $\delta_{c}=0$ is assumed per Remark 1) . Using the relation (9), the matrix $K_{\tau}$ is guaranteed to be positive definite if the primal regularization parameter $\delta_{w}$ is adequately large. As such, the parameter $\delta_{w}$ is chosen dynamically using the inertia information of the system in ($K_{1}$). Therefore, $K_{\tau}$ can be factorized with a Cholesky decomposition, satisfying the key requirement of stable pivoting for implementation on the GPU. The relaxation introduces an error in the final solution. Fortunately, the error is of the same order as the solver tolerance, thus it does not significantly deteriorate the solution quality for small $\varepsilon_{tol}$.
While this method can be implemented with minor modifications in the optimization solver, the presence of tight inequalities in (14) causes severe ill-conditioning throughout the IPM iterations. Therefore, using an accurate iterative refinement algorithm is necessary to achieve reliable convergence behavior.
@@ 3.3 Discussion
We have introduced two algorithms to solve KKT systems on the GPU. Unlike classical implementations, the two methods do not require computing a sparse $\mathrm{LBL^{T}}$ factorization of the KKT system and instead utilize alternate reformulations based on the condensed KKT system ($K_{1}$). Both strategies rely on a Cholesky factorization: HyKKT factors a positive definite matrix $K_{\gamma}$ obtained using an Augmented Lagrangian strategy, whereas LiftedKKT factors a positive definite matrix $K_{\tau}$ after employing an equality relaxation strategy. We will demonstrate in the next section that the ill-conditioned matrices $K_{\gamma}$ and $K_{\tau}$ possess a specific structure that limits the loss of accuracy in IPM.
@@ 4 Conditioning of the condensed KKT system
The relative errors are bounded above by a term depending on the conditioning $\kappa_{2}(M)$. Hence, it is legitimate to investigate the impact of the ill-conditioning when solving the condensed system ($K_{1}$) with LiftedKKT or with HyKKT. We will see that we can tighten the bounds in (16) by exploiting the structured ill-conditioning of the condensed matrix $K$. We base our analysis on wright1998ill , where the author has put a particular emphasis on the condensed KKT system ($K_{1}$) without equality constraints. We generalize their results to the matrix $K_{\gamma}$, which incorporates both equality and inequality constraints. The results extend directly to $K_{\tau}$ (by letting the number of equalities to be zero).
@@ 4.1 Centrality conditions
We start the discussion by recalling important results about the iterates of the interior-point algorithm. For $p:=(x,s,y,z)$, we denote by $(p,v)$ the current primal-dual iterate, and $(p^{\star},v^{\star})$ a solution of the KKT conditions (4). We denote $\mathcal{B}=\mathcal{B}(x^{\star})$ the active-set at the optimal solution $x^{\star}$, and $\mathcal{N}=\mathcal{N}(x^{\star})$ the inactive set. In this section, we are interested in the local convergence behavior of the primal-dual iterate, and we suppose $(p,v)$ is close enough to the solution $(p^{\star},v^{\star})$.
@@
Let $(p^{\star},v^{\star})$ be a primal-dual solution satisfying the KKT conditions (4). Let the following hold:
@@
[…]
@@
for some constants $C>0$ and $\alpha\in(0,1)$. Conditions (19b) ensure that the products $s_{i}v_{i}$ are not too disparate in the diagonal term $D_{s}$. This condition is satisfied (even if rather loosely) in the solver Ipopt (see (wachter2006implementation, , Equation (16))).
@@ 4.2.1 Invariant subspaces in $K_{\gamma}$
We recall that $m_{e}$ is the number of equality constraints, and define $\ell:=m_{e}+m_{a}$.
We express the structured ill-conditioning of $K_{\gamma}$ by modifying the approach outlined in (wright1998ill, , Theorem 3.2) to account for the additional term $\gamma G^{\top}G$ arising from the equality constraints. We show that the matrix $K_{\gamma}$ has two invariant subspaces (in the sense defined in (stewart1990matrix, , Chapter 5)), associated respectively to the range of the transposed active Jacobian (large space) and to the null space of the active Jacobian (small space).
@@
    In late IPM iterations, the active slacks are converging to 0. We observe that if $\frac{1}{\Xi}\leq\gamma\leq\frac{1}{s_{min}}$ the parameter $\gamma$ does not increase the ill-conditioning of the condensed matrix $K_{\gamma}$.
@@ 4.2.3 Numerical solution of the condensed system
Condition (a) implies both $s_{min}=\Theta(\Xi)$ and $s_{max}=\Theta(\Xi)$ (Proposition 1). Condition (b) supposes in addition $\gamma=\Theta(\Xi^{-1})$, making the matrix $\Sigma_{L}$ well-conditioned with $\underline{\sigma}=\Theta(\Xi^{-1})$, $\overline{\sigma}=\Theta(\Xi^{-1})$ and $\overline{\sigma}/\underline{\sigma}=\Theta(1)$. This is in line with the recommendation in golub2003solving , where the authors recommend chosing $\gamma=\frac{\|K\|}{\|G\|^{2}}$ (as here the condition (a) implies that $\|K\|=\Theta(\Xi)$). Condition (c) ensures that $\kappa_{2}(K_{\gamma})=\Theta(\overline{\sigma})$ satisfies $\kappa_{2}(K_{\gamma})\mathbf{u}\ll 1$ (implying the Cholesky factorization runs to completion). Condition (d) tells us that the perturbation caused by the Cholesky factorization is $\Delta_{s}K_{\gamma}=O(\mathbf{u}\|K_{\gamma}\|)$. As (30) implies $\|K_{\gamma}\|=\Theta(\Xi^{-1})$, we can incorporate $\Delta_{s}K_{\gamma}$ in the perturbation $\Delta K_{\gamma}$ given in Proposition 2.
@@ 4.3 Solution of the condensed KKT system
$\left\{\begin{aligned} \bar{r}_{1}&=-\widehat{r}_{1}+\widehat{H}^{\top}\big{(}\widehat{D}_{s}\widehat{r}_{4}-\widehat{r}_{2}\big{)}\;,\\ \bar{r}_{2}&=-\widehat{r}_{3}\;.\end{aligned}\right$. (40)
@@ 4.3.3 Summary
Most of the error arises in the descent w.r.t. the active bound’s multipliers $\widehat{d}_{z,\mathcal{B}}$. The impact remains limited if we have only few active inequality constraints.
5 Implementation
All our implementations use the Julia language bezanson-edelman-karpinski-shah-2017 . We utilized our local workstation to generate the results on the CPU, which is equipped with an AMD EPYC 7443 processor (3.2GHz) with 24 processors. For the GPU results, we used an NVIDIA A100 GPU (with CUDA 12.3) on the Polaris testbed at Argonne National Laboratory  [^1] https://www.alcf.anl.gov/polaris.
@@ 5 Implementation
We have implemented the two condensed-space methods in our nonlinear IPM solver MadNLP shin2021graph . This implementation utilizes the abstraction AbstractKKTSystem in MadNLP to represent the various formulations of the KKT linear systems. MadNLP can push most of the IPM algorithm to the GPU. In particular, any operation that involves the manipulation of array entries is performed by GPU kernels without transferring data to the host memory. We refer to shin2023accelerating for a detailed description of the GPU implementation in MadNLP.
@@
For fast evaluation of the derivatives on the GPU, we use the ExaModels.jl modeling tool shin2023accelerating to instantiate the evaluation of the nonlinear programs on the GPU. ExaModels.jl harnesses the sparsity structure and provides custom derivative kernels for repetitive algebraic subexpressions of the constraints and objective functions to compute first and second-order derivatives on the GPU in parallel bischof1991exploiting ; enzyme2021 . This approach caters to the SIMD architecture of the GPU by assigning each expression to multiple threads responsible for computing derivatives for different values. ExaModels.jl is efficient at evaluating the problem’s derivatives and offers a speedup often greater than 100 compared to standard tools like AMPL or JuMP (see Figure 1). As a result, the time spent in the evaluation of the model becomes negligible compared to the time spent in the linear solver. ExaModels.jl compiles the derivative evaluation kernels specifically for the given problem, whereas JuMP and AMPL depend on a precompiled interpreter to evaluate the model’s derivatives. Conversely, ExaModels.jl necessitates compilation time to create the derivative evaluation kernels, which can become substantial for problems involving a large number of heterogeneous constraints. Within this paper, we have always excluded the compilation time from the reported timings.
Figure 1: Evolution of the time spent evaluating the problem’s derivatives (gradient, Jacobian, Hessian) as we increase the problem’s dimension nh. In this benchmark, we compare four different instances from the COPS benchmark dolan2002benchmarking : rocket, robot, pinene, and dirichlet.
@@
We solve the KKT systems assembled within MadNLP using various sparse linear solvers, chosen based on the KKT formulation $($$K_{1}$), ($K_{2}$), ($K_{3}$)) and the device (CPU, GPU) being utilized. We utilize the following solvers:
  • HSL MA27: Implements the multifrontal $\mathrm{LBL^{T}}$ factorization on the CPU duff1983multifrontal .
  • HSL MA57: Implements the multifrontal $\mathrm{LBL^{T}}$ factorization on the CPU duff1983multifrontal , using multithreaded BLAS kernels in the elimination algorithm.
  • HSL MA86: Implements a fine-grained multicore parallel supernodal $\mathrm{LBL^{T}}$ factorization on the CPU.
  • Panua Pardiso: Implements shared-memory parallel supernodal Cholesky, $\mathrm{LDL^{T}}$ and $\mathrm{LBL^{T}}$ factorizations on the CPU.
  • CHOLMOD: Implements the Cholesky and $\mathrm{LDL^{T}}$ factorizations on the CPU (using the AMD ordering amestoy-david-duff-2004 by default). It factorizes the condensed matrices $K_{\gamma}$ and $K_{\tau}$ appearing respectively in (11) and in (15).
  • cuDSS: Implements Cholesky, $\mathrm{LDL^{T}}$ and $\mathrm{LU}$ decompositions on an NVIDIA GPU. We use the $\mathrm{LDL^{T}}$ factorization to factorize the ill-conditioned condensed matrices $K_{\gamma}$ on the GPU, (here, we have observed that $\mathrm{LDL^{T}}$ is more robust than the Cholesky factorization).
@@
6 Numerical results
First, we assess in §6.1 the performance of the two hybrid solvers, LiftedKKT and HyKKT. The GPU implementation is compared to state-of-the-art solvers running on the CPU. Then, we present in §6.2 the results reported on the PGLIB OPF benchmark, complemented in §6.3 by the CUTEst benchmark.
6.1 Performance analysis of HyKKT and LiftedKKT on a large-scale instance
We evaluate the performance of each KKT solver on a large-scale OPF instance, taken from the PGLIB benchmark babaeinejadsarookolaee2019power : 78484epigrids. Our formulation with ExaModels has a total of 674,562 variables, 661,017 equality constraints, and 378,045 inequality constraints. Our previous work has pointed out that as soon as the OPF model is evaluated on the GPU using ExaModels, the time spent in AD becomes negligible, and the KKT solver becomes the bottleneck in the numerical implementation shin2023accelerating .
6.1.1 Performance comparison of linear solvers on CPUs and GPUs
We evaluate the individual performance of the cuDSS solver when factorizing the matrix $K_{\gamma}$ at the first IPM iteration (here with $\gamma=10^{7}$). We compare the times to perform the symbolic analysis, the factorization, and the triangular solves with those reported in CHOLMOD (running single-threaded), HSL MA86, and Panua Pardiso with sparse Cholesky. The CPU implementations achieved their highest efficiency when configured with 8 threads. Therefore, in the remainder of this paper, we will consistently present the performance of the linear solvers operating in parallel using 8 threads.
The results are displayed in Table 1. We benchmark in parallel the three decompositions implemented in cuDSS (Cholesky, $\mathrm{LDL^{T}}$, $\mathrm{LU}$). We observe that the analysis phase is four times slower for cuDSS compared to CHOLMOD. The ordering happening within the symbolic analysis is performed on CPU, so it cannot benefit from the GPU parallelism nvidiaNVIDIACuDSSPreview . Fortunately, the analysis can be performed only once, and the symbolic factorization can be reused if the optimization problem is re-solved with the same sparsity structure. The numerical factorization is about two times faster in cuDSS than in HSL MA86 , with a time almost independent of the factorization being used.
analysis (s) factorization (s) backsolve (s)
CHOLMOD $1.37\times 10^{0}$ $7.44\times 10^{-1}$ $5.20\times 10^{-2}$
PARDISO $4.52\times 10^{0}$ $2.94\times 10^{-1}$ $8.94\times 10^{-2}$
MA86 $3.88\times 10^{0}$ $1.38\times 10^{-1}$ $5.50\times 10^{-2}$
cuDSS-cholesky $3.63\times 10^{0}$ $4.05\times 10^{-2}$ $7.22\times 10^{-3}$
cuDSS-ldl $3.57\times 10^{0}$ $4.21\times 10^{-2}$ $6.35\times 10^{-3}$
cuDSS-lu $3.55\times 10^{0}$ $6.28\times 10^{-2}$ $6.42\times 10^{-3}$
Table 1: Comparing the performance of cuDSS with CHOLMOD , Pardiso, and HSL MA86 (both running in parallel using 8 threads) . The column for “analysis” displays the time spent in the symbolic analysis (computation of a reordering to minimize fill-in and construction of the elimination tree). The columns for “factorization” and “backsolves” display the time spent in the numerical factorization and in the ensuing triangular solve, respectively. The matrix $K_{\gamma}$ is symmetric positive definite, with a size $n=674,562$. The matrix is extremely sparse, with only $7,342,680$ non-zero entries (0.0020.002%).